# Lattice Boltzmann methods for soft flowing matter

### v1

@@ IV Recent Lattice Boltzmann HPC implementations
In recent years, much work has been directed to developing advanced computational tools for the simulations of complex fluids on high performance computing (HPC) architectures. Indeed, the investigation of the flowing properties of confined soft materials often poses major multiscale challenges, due to the need to describe multi-component systems in centimeter-sized devices while retaining the physics of near-contact interactions. In this respect, standard LB approaches suffer a series of issues, the first being the relatively low arithmetic intensity, which varies between 1 and 5 Flops/Byte ratios depending on the specific implementation. Consequently, the LB method is often recognized as a memory-bound problem, as its efficiency is constrained by the memory access time on computing systems Succi et al. (2019). This problem is due to the necessity to directly store probabilistic distributions involving a set of approximately $b\sim 20-30$ discrete populations $f_{i}$ (where $i=0,...,b$ for each lattice point), clearly outnumbering the number of related hydrodynamic fields, which are a single scalar density $\rho$, the flow velocity $u_{\alpha}$ ($\alpha=1,3$), and momentum-flux tensor $\Pi_{\alpha\beta}$ (a $3\times 3$ symmetric tensor), amounting to 10 independent fields in total. Therefore, the LB method demands roughly double the memory compared to a corresponding computational fluid dynamics method based on Navier-Stokes equations. This redundancy buys several major computational advantages, primarily the fact that streaming is exact and diffusion is emergent (no need for second-order spatial derivatives to describe dissipative effects), which proved invaluable assets in achieving outstanding parallel efficiency across virtually any HPC platforms, also in the presence of real-world complex geometries Succi (2018). These extra-memory requirements put a premium on strategies aimed at minimizing the cost of data access in massively parallel LB codes.
@@
A number of other approaches focus on reducing memory occupancy without significantly compromising the simulation quality. They can be categorized into three groups: the first one centers on the algorithmic implementation of the streaming step, the second one leverages diverse numerical representations to rounding in floating point arithmetic, and the third one avoids the direct storage of the probabilistic information by exploiting a reconstruction procedure of distributions from the hydrodynamics fields via Hermite projection.
@@
Second group.
The second strategy employs a single precision (instead of a double one) or a mixed single precision/half-precision representation which reduces the total memory usage by a factor of two without compromising the simulation quality Lehmann et al. (2022); Gray and Boek (2016). This approach was initially introduced by Skordos Skordos (1993), who showed that if the distribution functions are adjusted by negating the equilibrium zero-velocity distribution function, the probabilistic populations have the potential to maintain more significant bits during the execution of floating-point operations, thus increasing the computational precision. This idea was further developed in Ref.Gray and Boek (2016), where the dependence on the velocity of the accuracy of LB calculations, affecting the previous approach, was removed. More specifically, this method requires a modification of distribution functions and moments as follows. Using the symbols (float32) and (float64) to indicate the cast operators forcing a data type to be converted into single (32 bits/4 bytes) and double (64 bits/8 bytes) floating point precision, and writing
@@
$[\delta f_{i}({\bf x}+{\bf c}_{i}\Delta t,t+\Delta t)]_{\text{32bit}}=(\text{float32})\biggl{\{}\omega\delta f_{i}^{eq}+$
$(1-\omega)(\text{float64})[\delta f_{i}({\bf x},t)]_{\text{32bit}}\biggr{\}},$ (73)
@@
$\delta f_{i}^{eq}=f_{i}^{eq}(\rho,{\bf u})-f_{i}^{eq}(\rho_{0},{\bf u}_{0})=w_{i}\Big{[}\frac{{\bf c}_{i}\cdot{\bf u}}{c_{s}^{2}}+\frac{({\bf c}_{i}{\bf c}_{i}-c_{s}^{2}\mathbf{I}):{\bf u}{\bf u}}{2c_{s}^{4}}\Big{]}+w_{i}(\rho-\rho_{0}),$ (74)
@@
Note that the distributions $\delta f_{i}(\vec{x},t)$ are centered around zero and shifted by $f_{i}^{eq}(\rho_{0},\vec{u}_{0})$, which means that the accuracy of the summation of small differences is not limited by the order of magnitude of the density $\rho_{0}$. Further, the summation can be performed with double precision accuracy. Recently, Lehmann et al. Lehmann et al. (2022) implemented the same strategy exploiting a mixed precision approach (FP-32/FP-16), where the floating-point operations are carried out in single precision, while the extra population terms are saved in half-precision (16 bits/2 bytes) $[\delta f_{i}(\vec{x},t)]_{\text{16bit}}$. In particular, they introduced a customized half-precision number format (FP16C) that halved truncation error compared to the standard IEEE-754 half-precision floating-point format FP16 in LBM applications Lehmann et al. (2022). However, benchmarks on the Karman vortex street in two dimensions have shown the presence of numerical noise in the third decimal digits, where the vorticity is compared between the mixed precision platforms FP64/32 and FP32/16C (see Fig.11). In concrete numbers, it is possible to quantify the LB performance in terms of a billion lattice sites per second (GLUPS), namely the number of updated lattice nodes per second reported in billions. For instance, Lehmann et al. Lehmann et al. (2022) measured the performance equal to 8.5 GLUPS on a single GPU NVIDIA A100 PCIe with 40GB RAM implementing the AA-Pattern in-place streaming scheme for a single component BGK LB model in a single precision floating point precision for storage and floating point operations and achieving a performance peak of 16 GLUPS by adopting the mixed precision approach (FP-32/FP-16).
@@
Third group.
The third strategy for managing memory usage is grounded on the ideas of Ladd and Verberg Ladd and Verberg (2001), in which a memory reduction could be achieved by solely storing hydrodynamic fields and their gradients. This concept hinges on the fact that the probabilistic data needed to execute any LB method can be dynamically reconstructed using the available hydrodynamic quantities, thus eliminating the need for population storage. This is especially beneficial for multi-component and multi-species applications transported by a common flow field in which a single hydrodynamic field (i.e. the density) is needed rather than a full kinetic representation with $O(30)$ populations. A similar logic applies to flows that are far from equilibrium or involve relativistic hydrodynamics, which require higher-order lattices, sometimes involving hundreds of discrete populations per species. All methods based on this third approach typically rely on a moment-based representation of the LB, bypassing the direct storage of the probabilistic populations along each direction of the lattice. Among these, noteworthy implementations include the recently introduced lightweight LB (LLB) and thread-safe LB (TSLB) methods Tiribocchi et al. (2023c); Montessori et al. (2023, 2024), which are based on the reconstruction of probabilistic distributions via Hermite projection.
@@
$f_{i}({\bf x}+{\bf c}_{i}\Delta t,t+\Delta t)=f^{eq}_{i}\biggl{(}\rho({\bf x},t),{\bf u}({\bf x},t)\biggr{)}+$
$\left(1-\omega\right)f^{neq}_{i}\biggl{(}\Pi^{neq}_{\alpha\beta}({\bf x},t)\biggr{)},$ (81)
@@
This procedure has been used to implement LLB and TSLB models described in Refs.Tiribocchi et al. (2023c); Montessori et al. (2023) for single and two-component fluids. In Fig.12 we show, for example, an off-axis collision between equal-size fluid droplets simulated using the LLB method (a)-(e) presented in Ref.Tiribocchi et al. (2023c) and the standard LB method (f)-(j). In both cases, the relative impact speed is fixed at 0.5 in lattice units, the Reynolds and Capillary numbers for the droplet (red fluid) are ReR = 450 and CaR $\simeq 1.6$, and ReB = 50 for the surrounding medium (blue fluid). The time sequence in Fig.12 clearly shows that both methods result in essentially similar dynamic behaviors.
@@
Figure 12: A sequence of an off-axis collision between two fluid droplets simulated using the LLB scheme without populations (a)-(e) and a standard LB scheme with populations (f)-(j). The centers of mass are initially positioned at a distance where the droplets’ mutual interaction is minimal. The droplets, starting as perfect spheres [(a)-(f)], acquire a bullet-like shape as they collide (b), (c) and (g), (h). They then separate (d)-(i) and regain a spherical shape once sufficiently far apart (e)-(j). There is no significant difference between the two methods. The color map shows the density field values, which range from 0 to 1. Adapted with permission from Ref. Tiribocchi et al. (2023c).
@@
We conclude this section by noting that, over the last two decades, several examples of moment-based models have been published, such as the CPU-based one presented in Ref.Argentini et al. (2004) and more sophisticated releases implemented on GPU computers Ferrari et al. (2023); Gounley et al. (2021); Vardhan et al. (2019), where ad-hoc memory layouts are adopted for the hydrodynamics array to capitalize the full memory bandwidth of multi-GPU clusters Ferrari et al. (2023). These techniques have been shown to enhance data continuity, reducing the impact of memory-bound dependence on the implementation Bader (2012) with benefits in terms of computational performance. As a concrete example, Ferrari et al. Ferrari et al. (2023) measured a peak of 10.5 GLUPS on a single GPU NVIDIA A100 PCIe with 40GB RAM adopting a z-curve memory layout for a single component regularized-LB model with an increase of about 20% compared to what observed by Lehmann et al. Lehmann et al. (2022) for a single component BGK-LB model at the same floating point precision.
@@ V Selected applications
In this section, we focus on specific applications which have mostly benefited from the LB approaches and where this method is currently playing a game-changing role, meaning by this that the computational modeling of these applications would have been more demanding, if possible at all, than other methods. We refer primarily to a variety of droplet motions in microfluidic devices, the rheology of confined dense emulsions and other complex phenomena which may lay the ground for a new class of droplet-based materials. More specifically, we identify the following five instances where, we believe, the LB method has shown a major impact: soft flowing crystals, soft granular media, dense emulsions under thermal flows, hierarchical multiple emulsions, active gel droplets in highly confined environments (such as pore-sized constrictions) and extreme flow simulations modeling macroscopic biological entities, such as deep-sea sponges.

### v2

@@ IV Recent Lattice Boltzmann HPC implementations
In recent years, much work has been directed to developing advanced computational tools for the simulations of complex fluids on high performance computing (HPC) architectures. Indeed, the investigation of the flowing properties of confined soft materials often poses major multiscale challenges, due to the need to describe multi-component systems in centimeter-sized devices while retaining the physics of near-contact interactions. In this respect, standard LB approaches suffer a series of issues, the first being the relatively low arithmetic intensity, which varies between 1 and 5 Flops/Byte ratios depending on the specific implementation. Consequently, the LB method is often recognized as a memory-bound problem, as its efficiency is constrained by the memory access time on computing systems Succi et al. (2019). This problem is due to the necessity to directly store probabilistic distributions involving a set of approximately $b\sim 20-30$ discrete populations $f_{i}$ (where $i=0,...,b$ for each lattice point), clearly outnumbering the number of related hydrodynamic fields, which are a single scalar density $\rho$, the flow velocity $u_{\alpha}$ ($\alpha=1,3$), and momentum-flux tensor $\Pi_{\alpha\beta}$ (a $3\times 3$ symmetric tensor for an isotropic binary fluid), amounting to 10 independent fields in total. Therefore, the LB method demands roughly double the memory compared to a corresponding computational fluid dynamics method based on Navier-Stokes equations. This redundancy buys several major computational advantages, primarily the fact that streaming is exact and diffusion is emergent (no need for second-order spatial derivatives to describe dissipative effects), which proved invaluable assets in achieving outstanding parallel efficiency across virtually any HPC platforms, also in the presence of real-world complex geometries Succi (2018). These extra-memory requirements put a premium on strategies aimed at minimizing the cost of data access in massively parallel LB codes.
@@
A number of other approaches focus on reducing memory occupancy without significantly compromising the simulation quality. They can be categorized into three groups: the first one centers on the algorithmic implementation of the streaming step, the second one leverages diverse numerical representations to rounding in floating-point arithmetic, and the third one avoids the direct storage of the probabilistic information by exploiting a reconstruction procedure of distributions from the hydrodynamics fields via Hermite projection. Below we describe some features of each of these methods and then we compare the computational performance of a recent version of the LB pertaining to the third group with Navier-Stokes solvers.
@@ IV.1 First group
IV.2 Second group
The second strategy employs a single precision (instead of a double one) or a mixed single precision/half-precision representation which reduces the total memory usage by a factor of two without compromising the simulation quality Lehmann et al. (2022); Gray and Boek (2016). This approach was initially introduced by Skordos Skordos (1993), who showed that if the distribution functions are adjusted by negating the equilibrium zero-velocity distribution function, the probabilistic populations have the potential to maintain more significant bits during the execution of floating-point operations, thus increasing the computational precision. This idea was further developed in Ref.Gray and Boek (2016), where the dependence on the velocity of the accuracy of LB calculations, affecting the previous approach, was removed. More specifically, this method requires a modification of distribution functions and moments as follows. Using the symbols (float32) and (float64) to indicate the cast operators forcing a data type to be converted into single (32 bits/4 bytes) and double (64 bits/8 bytes) floating-point precision, and writing
@@ IV.2 Second group
$[\delta f_{i}({\bf x}+{\bf c}_{i}\Delta t,t+\Delta t)]_{\text{32bit}}=(\text{float32})\biggl{\{}\omega\;\delta f_{i}^{eq}$
$+(1-\omega)(\text{float64})[\delta f_{i}({\bf x},t)]_{\text{32bit}}\biggr{\}},$ (73)
@@
$\delta f_{i}^{eq}=f_{i}^{eq}(\rho,{\bf u})-f_{i}^{eq}(\rho_{0},{\bf u}_{0})=$
$w_{i}\Big{[}\frac{{\bf c}_{i}\cdot{\bf u}}{c_{s}^{2}}+\frac{({\bf c}_{i}{\bf c}_{i}-c_{s}^{2}\mathbf{I}):{\bf u}{\bf u}}{2c_{s}^{4}}\Big{]}+w_{i}(\rho-\rho_{0}),$ (74)
@@
Note that the distributions $\delta f_{i}(\vec{x},t)$ are centered around zero and shifted by $f_{i}^{eq}(\rho_{0},{\bf u}_{0})$, which means that the accuracy of the summation of small differences is not limited by the order of magnitude of the density $\rho_{0}$. Further, the summation can be performed with double precision accuracy.
Recently, Lehmann et al. Lehmann et al. (2022) implemented the same strategy exploiting a mixed precision approach (FP-32/FP-16), where the floating-point operations are carried out in single precision, while the extra population terms are saved in half-precision (16 bits/2 bytes) $[\delta f_{i}(\vec{x},t)]_{\text{16bit}}$. In particular, they introduced a customized half-precision number format (FP16C) that halved truncation error compared to the standard IEEE-754 half-precision floating-point format FP16 in LBM applications Lehmann et al. (2022). However, benchmarks on the Karman vortex street in two dimensions have shown the presence of numerical noise in the third decimal digits, where the vorticity is compared between the mixed precision platforms FP64/32 and FP32/16C (see Fig.11).
In concrete numbers, it is possible to quantify the LB performance in terms of billion lattice sites per second (GLUPS), namely the number of updated lattice nodes per second reported in billions. For instance, Lehmann et al. Lehmann et al. (2022) measured a performance equal to 8.5 GLUPS on a single GPU NVIDIA A100 PCIe with 40GB RAM implementing the AA-Pattern in-place streaming scheme for a single component BGK LB model in a single precision floating-point, and a performance peak of 16 GLUPS by adopting the mixed precision approach (FP-32/FP-16). A list of benchmarks carried out using their code (named FluidX3d, freely available on GitHub Lehmann (2024)) on several GPU and CPU-based devices are reported in Table 1. They were performed using a standard BGK operator with a D3Q19 scheme on a three-dimensional cubic box of 256 nodes per side, for different mixed floating-point representations.
@@
CPU hardware FP32/FP32 FP32/FP16 FP32/FP16C
2x Intel Xeon CPU Max 9480 2.037 1.520 1.464
2x Intel Xeon Platinum 8380 1.410 1.159 1.298
Intel Core i7-13700K 0.504 0.398 0.424
Intel Core i7-9700 0.103 0.062 0.095
GPU hardware FP32/FP32 FP32/FP16 FP32/FP16C
Nvidia H100 NVL 20.018 32.613 17.605
Nvidia A100 PCIe 40GB 8.526 16.035 11.088
Nvidia V100 PCIe 16GB 5.128 10.325 7.683
Nvidia GeForce RTX 4090 5.624 11.091 11.496
Table 1: Benchmarks performed by a standard BGK operator with the D3Q19 scheme on a three-dimensional cubic box of 256 nodes per side for different mixed floating point-representations. The results are reported in GLUPS for several Intel CPU and Nvidia GPU architectures. The table is reproduced with permission from Ref. Lehmann (2024).
IV.3 Third group
The third strategy for managing memory usage is grounded on the ideas of Ladd and Verberg Ladd and Verberg (2001), in which a memory reduction could be achieved by solely storing hydrodynamic fields and their gradients. This concept hinges on the fact that the probabilistic data needed to execute any LB method can be dynamically reconstructed using the available hydrodynamic quantities, thus eliminating the need for population storage. This is especially beneficial for multi-component and multi-species applications transported by a common flow field in which a single hydrodynamic field (i.e., the density) is needed rather than a full kinetic representation with $O(30)$ populations. A similar logic applies to flows that are far from equilibrium or involve relativistic hydrodynamics, which require higher-order lattices, sometimes involving hundreds of discrete populations per species. All methods based on this third approach typically rely on a moment-based representation of the LB, bypassing the direct storage of the probabilistic populations along each direction of the lattice. Among these, noteworthy implementations include the recently introduced lightweight LB (LLB) and thread-safe LB (TSLB) methods Tiribocchi et al. (2023c); Montessori et al. (2023, 2024), which are based on the reconstruction of probabilistic distributions via Hermite projection.
@@ IV.3 Third group
$f_{i}({\bf x}+{\bf c}_{i}\Delta t,t+\Delta t)$ $=f^{eq}_{i}\biggl{(}\rho({\bf x},t),{\bf u}({\bf x},t)\biggr{)}$ (81)
$+\left(1-\omega\right)f^{neq}_{i}\biggl{(}\Pi^{neq}_{\alpha\beta}({\bf x},t)\biggr{)},$
@@
This procedure has been used to implement LLB and TSLB models described in Refs.Tiribocchi et al. (2023c); Montessori et al. (2023) for single and two-component fluids. In Fig.12 we show, for example, an off-axis collision between equal-size fluid droplets in a cubic box, simulated using the LLB method presented in Ref.Tiribocchi et al. (2023c) (Fig.12a-e) and the standard LB method (Fig.12f-j). In both cases, the relative impact speed is fixed at 0.5 in lattice units, the Reynolds and Capillary numbers for the droplet (red fluid) are ReR = 450 and CaR $\simeq 1.6$, and ReB = 50 for the surrounding medium (blue fluid). The time sequence of the collision plus the plots of density and relative error in Fig.12 clearly show that both methods result in essentially similar dynamic behaviors.
@@
Figure 12: A sequence of an off-axis collision between two fluid droplets simulated using the LLB scheme without populations (a)-(e) and a standard LB scheme with populations (f)-(j). The centers of mass are initially positioned at a distance where the droplets’ mutual interaction is minimal. The droplets, starting as perfect spheres [(a)-(f)], acquire a bullet-like shape as they collide (b), (c) and (g), (h). They then separate (d)-(i) and regain a spherical shape once sufficiently far apart (e)-(j). There is no significant difference between the two methods. The color map shows the density field values, which range from 0 to 1. k) Time evolution of the fluid density $\rho$ computed using a standard LB (POPS) and the lightweight LB (NOPOPS), where delimited-⟨$\langle$..$\rangle$ … ⟩ is a spatial average on lattice sites with $\rho>0.5$ (i.e. within the droplets, red region). The interaction time between the droplets is $10^{2}\lesssim t\lesssim 15\times 10^{2}$. l) Time evolution of the relative error $|\rho_{N}-\rho_{P}|/\rho_{N_{MAX}}$ at three different lattice sites, i.e. near the droplet interface $(x=32,y=64,z=64)$, within the droplet $(x=45,y=50,z=64)$ and far from the droplets $(x=128,y=128,z=128)$. Here, $\rho_{N}$ is the density computed using the LLB and $\rho_{P}$ is the one from the standard LB. The figure is adapted with permission from Ref. Tiribocchi et al. (2023c).
@@
We conclude noting that, over the last two decades, several examples of moment-based models have been published, such as the CPU-based one presented in Ref.Argentini et al. (2004) and more sophisticated releases implemented on GPU computers Ferrari et al. (2023); Gounley et al. (2021); Vardhan et al. (2019), where ad-hoc memory layouts are adopted for the hydrodynamics array to capitalize the full memory bandwidth of multi-GPU clusters Ferrari et al. (2023). These techniques have been shown to enhance data continuity, reducing the impact of memory-bound dependence on the implementation Bader (2012) with benefits in terms of computational performance. As a concrete example, Ferrari et al. Ferrari et al. (2023) measured a peak of 10.5 GLUPS on a single GPU NVIDIA A100 PCIe with 40GB RAM adopting a z-curve memory layout for a single component regularized-LB model with an increase of about 20% compared to what observed by Lehmann et al. Lehmann et al. (2022) for a single component BGK-LB model at the same floating-point precision.
IV.4 Performance comparison between lattice Boltzmann and Navier-Stokes equation solvers
We close this section by providing a quantitative comparison, in terms of GLUPS, between LB methods and Navier-Stokes solvers. We begin by considering the declared performance of the commercial computational fluid dynamic code ANSYS Fluent Matsson (2023) based on the volume of fluid method, focusing on its GPU-accelerated version. For a single-component CFD simulation on a cubic grid of 126 lattice nodes per side (roughly 2 million nodes in total), the reported performance on a single GPU (NVidia A100) is 0.042 GLUPS, scaling up to 0.266 GLUPS when using 8 GPUs. In comparison, a recent version of the LB method Lehmann (2024), operating on a cubic grid of 256 lattice nodes per side (more than 16 million nodes in total), achieves approximately 8.5 GLUPS on a single GPU (see Table 1), roughly 200 times faster than the single-GPU performance of ANSYS code. It is worth noting that other non-commercial, GNU-licensed Navier-Stokes solvers, which are often optimized for the simulation of specific flow types (see Costa et al. (2021), for example), show better performance than ANSYS, typically up to an order of magnitude faster. However, even in these cases, LB methods still demonstrate higher performance when it comes to updating a single computational node.
Regarding massively parallel implementations, GNU-licensed Navier-Stokes solvers show high scalability on multi-GPU clusters. The scalability can be evaluated in terms of the speedup, defined as the ratio between the CPU wall-clock time of a code executed on a single core and the one of the code executed in parallel mode. For instance, the FluTAS solver, which uses finite difference methods (including a pseudo-spectral Poisson solver) to simulate multiphase flows with thermal effects, shows a speed-up of about 83, with performance ranging from 0.07 GLUPS on 8 GPUs NVidia A100 to almost 0.73 GLUPS on 128 GPUs devices on a grid of $1024\times 512\times 256$ lattice points Crialesi-Esposito et al. (2023). On the contrary, a recent implementation of the lattice Boltzmann, named LBcuda Bonaccorso et al. (2022), proved a speed-up equal to 34 on 64 GPUs NVidia A100 devices in a cubic box of linear size 512 (see Fig.13) and, more generally, a good scaling for increasing number of GPU cards (at least for a sufficiently large system size). It is finally important to clarify that this comparison does not account for the ability of standard CFD methods to work with unstructured or non-uniform meshes, which remains an area where traditional Navier-Stokes solvers have certain advantages.
Figure 13: Speed-up of the LBcuda code using different A100 GPUs (each with 80 GB of memory), tested on two cubic grids with side lengths of 256 and 512, respectively. The figure is adapted with permission from Ref.Bonaccorso et al. (2022).
@@ V Selected applications
In this section, we focus on specific applications which have mostly benefited from the LB approaches and where this method is currently playing a game-changing role, meaning by this that the computational modeling of these applications would have been more demanding, if possible at all, than other methods. We distinguish systems for which a validation against quantitative results has been successfully demonstrated, from simulations often inspired by experiments in contiguous fields of research and where the agreement is mostly built on qualitative grounds. Rather than being a specific limitation, this conveys the idea that the method, besides being capable of capturing some key features of the physics of highly complex systems, has the potential to predict behaviors not yet explored in the context of soft flowing materials. We refer primarily to a variety of droplet motions in microfluidic devices, the rheology of confined dense emulsions and other complex phenomena which may lay the ground for a new class of droplet-based materials. More specifically, we identify the following five instances where, we believe, the LB method has shown a major impact: soft flowing crystals, soft granular media, dense emulsions under thermal flows, hierarchical multiple emulsions, active gel droplets in highly confined environments (such as pore-sized constrictions) and extreme flow simulations modeling macroscopic biological entities, such as deep-sea sponges.